𝐴 = { 𝑧 ∈ β„‚:|𝑧| < 1 𝑣𝑒 |𝑧 βˆ’ 1/ 2 | > 1 /2 } βˆͺ { 1/ 2 } denote the set in the complex plane and

𝐴 β€² and πœ•π΄ = 𝐴̅ \ π΄π‘œ

Write the set.


1
Expert's answer
2020-08-10T17:56:43-0400

the set in the complex plane:



Write the set:

"\overline{A} = \{ z \in \Complex : |z| \ge 1; |z-\frac{1}{2}| \le \frac{1}{2} \} \setminus \{ \frac{1}{2} \}
\newline
\partial A = \{ z \in \Complex : |z| = 1; |z - \frac{1}{2}| = \frac{1}{2} \}"